OpenStudy (anonymous):

find arclength: x=arcsint y=ln square root(1-x^2) x is between 0 and 1/2

OpenStudy (amistre64):

\[\int_{sin(0)}^{sin(1/2)}\sqrt{(x')^2+(y')^2}dt\] maybe?

OpenStudy (amistre64):

x = arcsint defines x as an angle

OpenStudy (amistre64):

you sure you got this posted right?

OpenStudy (anonymous):

ya, oopsss. it should be y= ln(square root of( 1-(t^2))

OpenStudy (amistre64):

that does make a little more sense to me :)

OpenStudy (amistre64):

arcsin' = 1/sqrt(1-t^2) , if my memory serves me

OpenStudy (amistre64):

x = arcsint sinx = t cosx x' = 1 x' = 1/cosx = 1/sqrt(1-sin^2x) = 1/sqrt(1-t^2)

OpenStudy (amistre64):

y= ln(square root of( 1-(t^2)) y = 1/2 ln (1-t^2) 2y = ln(1-t^2) 2y' = -2t/(1-t^2) y' = -t/(1-t^2)

OpenStudy (amistre64):

squre those up; add them; and sqrt the results :)

OpenStudy (anonymous):

thanx